### v1

# Tazza: Shuffling Neural Network Parameters for Secure and Private Federated Learning

Abstract.
Federated learning enables decentralized model training without sharing raw data, preserving data privacy. However, its vulnerability towards critical security threats, such as gradient inversion and model poisoning by malicious clients, remain unresolved. Existing solutions often address these issues separately, sacrificing either system robustness or model accuracy. This work introduces Tazza, a secure and efficient federated learning framework that simultaneously addresses both challenges. By leveraging the permutation equivariance and invariance properties of neural networks via weight shuffling and shuffled model validation, Tazza enhances resilience against diverse poisoning attacks, while ensuring data confidentiality and high model accuracy. Comprehensive evaluations on various datasets and embedded platforms show that Tazza achieves robust defense with up to 6.7$\times$ improved computational efficiency compared to alternative schemes, without compromising performance.
copyright: none
1. Introduction
Federated learning is a transformative paradigm that facilitates training machine learning models in decentralized environments while preserving the privacy of data stored on distributed devices (McMahan et al., 2017). In federated learning, clients perform local model training over several epochs and share model updates, rather than raw data, with a central server. The server then aggregates these updates across clients over multiple rounds, collaboratively building a robust global model. This decentralized approach is particularly promising for mobile and embedded computing systems (Shin et al., 2024; Alam et al., 2023), enabling them to leverage vast amounts of locally collected data while delivering personalized models tailored to individual users (Sani et al., 2024).
Despite its decentralized and privacy-preserving design, federated learning still faces significant privacy and security threats. Unlike centralized training, where data resides on a single server, federated learning relies on distributed, heterogeneous client data, leaving it vulnerable to malicious model updates that compromise system integrity (Fang et al., 2020). Furthermore, honest-but-curious servers can analyze model updates to infer private client data, breaching confidentiality. These concerns are particularly critical in mobile and embedded systems, where sensitive data is abundant, and the software environment is vulnerable due to the ease of installing third-party apps. These threats are categorized into two primary attack vectors: integrity attacks, in which malicious clients undermine model robustness, and confidentiality attacks, where curious servers endanger data privacy.
Specifically, integrity attacks involve malicious clients disrupting federated learning by sharing tampered updates, such as injecting noise (Fang et al., 2020), training with incorrect labels (Biggio et al., 2012; Fung et al., 2020), or embedding adversarial patterns (Bagdasaryan et al., 2020). Without effective defense, these integrity attacks hinder global model convergence and performance. Existing solutions include filtering tampered updates using statistical priors (Yin et al., 2018) or employing robust aggregation (Blanchard et al., 2017; Cao et al., 2020). Furthermore, confidentiality attacks, where the server aims to infer client data from model updates, are typically countered by techniques like differential privacy (Abadi et al., 2016; Dwork et al., 2014) or gradient pruning (Zhu et al., 2019). However, such methods often compromise global model convergence due to altered updates (Huang et al., 2021). This paper aims to comprehensively address these challenges.
We emphasize that ensuring both integrity and confidentiality is critical in federated learning, particularly for mobile and embedded systems. With large-scale participation from distributed and often anonymous clients, it is unrealistic to assume that all participants act in good faith. Robust mechanisms are needed to defend against malicious clients that compromise the integrity of the system. Equally important is safeguarding sensitive data collected from mobile and embedded devices, such as health information (Park et al., 2020), daily activities (Park et al., 2021; Ouyang et al., 2024; Fu et al., 2024), or environmental metrics (Xie et al., 2023; Sun et al., 2020) to maintain user trust and meet privacy regulations.
However, addressing integrity and confidentiality simultaneously poses unique challenges. Confidentiality defenses, such as noise addition, obscure model updates to protect privacy but can inadvertently hinder the detection of malicious patterns for backdoor attacks or label poisoning. At the same time, integrity defenses identify and mitigate such threats but may conflict with confidentiality measures, complicating the design of robust systems that target both objectives.
This challenge is amplified by the non-i.i.d. nature of real-world data. The diversity of benign model updates makes it difficult to differentiate legitimate updates from malicious ones, especially when compounded by the inaccuracies introduced by confidentiality defenses. These trade-offs create significant barriers to ensuring security and privacy without degrading global model performance.
Despite the critical importance of these issues, few approaches simultaneously address integrity and confidentiality threats while preserving model performance. Tackling this intertwined challenge is imperative for real-world federated learning applications, where system reliability and user trust depend on effective, balanced defense mechanisms.
In this paper, we address the challenges of ensuring both integrity and confidentiality in federated learning by leveraging the permutation equivariant and invariant properties of neural networks (c.f., Sec 2.2). Permutation equivariance ensures that a function’s output changes consistently with the order of its inputs, while permutation invariance guarantees the output remains unchanged regardless of input order. These properties are fundamental to widely used neural network architectures, including multilayer perceptrons (MLPs), recurrent neural networks (RNNs), and Transformers.
Leveraging these properties, we propose Tazza, a novel approach that simultaneously ensures confidentiality and integrity in federated learning. Specifically, by shuffling neural network model parameters, Tazza utilizes the permutation equivariance property to obscure sensitive data, enhancing confidentiality. Meanwhile, permutation invariance ensures that the shuffled parameters produce consistent computational outcomes, thereby maintaining integrity. This approach effectively demonstrates how permutation-based techniques can address fundamental security issues in federated learning without sacrificing model performance.
Nevertheless, properly shuffling model parameters while preserving desired properties is challenging, as it requires a deep understanding of neural network operations. While many architectures exhibit permutation equivariant and invariant properties, the mechanisms enabling these properties vary across models. To address this, we present tailored algorithms for representative architectures, offering a systematic approach to harness these properties. Our methods demonstrate the practical potential of permutation-based techniques in federated learning.
Building on the aforementioned properties, Tazza introduces a novel weight shuffling module, designed to obscure the original model by rearranging its weights while maintaining permutation equivariant and invariant properties. This ensures that the shuffled model remains valid, enabling us to propose a shuffled validation process for integrity checks, confirming that model outputs remain accurate despite weight shuffling. Finally, we leverage the similarity metrics obtained from the shuffled validation process and cluster the models to distinguish and isolate malicious updates from impacting benign models.
By integrating these components, Tazza emerges as a secure, privacy-preserving, and accurate federated learning framework that effectively addresses both integrity and confidentiality challenges. Along with a strong mathematical foundation leveraging permutation properties, we demonstrate its practical efficacy through extensive evaluations on diverse datasets and embedded platforms. Specifically, Tazza maintains high model accuracy while mitigating both integrity and confidentiality attacks, with a computational speedup of up to 6.7$\times$ compared to alternative defense schemes. Our key contributions are summarized as follows:
  • Building on mathematical evidence, we propose an effective and efficient model-securing mechanism tailored for federated learning scenarios. This mechanism leverages the permutation equivariance and invariance properties of neural networks through two key components: Weight Shuffling and Shuffled Model Validation.
  • By seamlessly integrating Weight Shuffling and Shuffled Validation, we present Tazza, a privacy-preserving and secure federated learning framework specifically designed for mobile and embedded scenarios, achieving robust security without compromising model performance.
  • We conduct an extensive evaluation of Tazza across diverse datasets, model architectures, and attack scenarios. Our results demonstrate that Tazza effectively isolates malicious model updates from benign ones to counter integrity attacks, while its model parameter shuffling mechanism robustly mitigates confidentiality attacks.
2. Background and Related Works
Figure 1. Integrity and confidentiality attack threats in a federated learning scenario.
Figure 2. (a) Visualization of the permutation equivariance property of matrix multiplication. The output’s column order follows the order of the input columns. (b) Visualization of the permutation invariance property of the average function. Results are not affected by the altered order of input rows. Best viewed in color.
2.1. Security Threats in Federated Learning
We begin by discussing background information on the two core security threats in federated learning: integrity and confidentiality attacks.
Integrity Attacks. Malicious clients can undermine the integrity of the federated learning process by compromising the reliability of model weights. Figure 1 (a) illustrates how such attacks operate. For instance, in data poisoning attacks, adversaries manipulate local models with altered data-label pairs (e.g., label flipping) to degrade global performance. In model poisoning attacks, they tamper with model weights by injecting noise (noise-injection) or scaling specific parameters (scaling attack) to distort outcomes (Fang et al., 2020). Additionally, backdoor attacks involve training models to produce undesired outputs when triggered by specific input patterns (Bagdasaryan et al., 2020). Collectively, these adversarial activities are referred to as integrity attacks in this paper.
To mitigate integrity attacks, prior research has proposed secure model aggregation techniques leveraging statistical priors or similarity-based approaches. Methods like Median (Xie et al., 2018) and Trimmed-Mean (Yin et al., 2018) aggregation exclude malicious updates by focusing on statistical measures rather than naive averaging, assuming that malicious updates are statistical outliers. Alternatively, similarity-based techniques such as Krum (Blanchard et al., 2017) use Euclidean distance to select updates with minimal deviation across models, while FLTrust (Cao et al., 2020) calculates trust scores based on cosine similarity with a verified, clean model stored at the server. Despite their strengths, these methods often discard valuable information in the updates, leading to suboptimal global model performance.
Confidentiality Attacks. Another significant security threat in federated learning systems is confidentiality attacks, which aim to infer and reveal local training data at the federated learning server. Figure 1 (b) illustrates the workflow of these attacks. Zhu et al. introduced DLG, demonstrating that training data can be reconstructed by analyzing the differences between the initial model distributed by the server and the updated model parameters returned by the local clients (Zhu et al., 2019). Specifically, DLG employs an optimization-based approach using L-BFGS to identify data that generates gradients minimizing the difference between the local model and the updated parameters, as shown in Figure 1 (b). Building on DLG, Geiping et al. proposed the Inverting Gradient attack, which incorporates the cosine similarity of the gradient and the total variation of the reconstructed image (Geiping et al., 2020).
To counter confidentiality attacks, widely adopted approaches include applying differential privacy (Abadi et al., 2016), which adds controlled noise to model updates, and gradient pruning (Zhu et al., 2019), which suppresses sensitive information in gradients during local training. However, these defense mechanisms often degrade model performance, as they manipulate original updates in ways that hinder convergence and accuracy (Huang et al., 2021).
2.2. Permutation Equivariance/Invariance
We now explore the concepts of permutation equivariance and permutation invariance, which are fundamental to the approach presented in this work.
Permutation Equivariance refers to the property where the order of a function’s output aligns with the order of its input. Figure 2 (a) illustrates this concept using matrix multiplication, where $R$ represents a column-reversed matrix. Specifically, the figure demonstrates that when the column order of the input matrix $X$ is reversed ($X\rightarrow X^{R}$), the resulting matrix multiplication outputs, $WX$ (top) and $WX^{R}$ (bottom), differ only in the column order, which corresponds to the input’s order.
Formally, if a function $f(\cdot)$ satisfies permutation equivariance, for a given input $X=\{x_{1},x_{2},\dots,x_{N}\}$ and a permutation function $\pi(i):\{1,2,\ldots,N\}\to\{1,2,\ldots,N\}$, where $\pi(i)$ is a bijective mapping, the lemma $\pi(f(X))=f(\pi(X))$ holds. Applying this to the case of matrix multiplication, the function $f(X)=WX$ satisfies permutation equivariance with respect to the column order. Consequently, $WX^{R}$ is identical to $(WX)^{R}$.
Permutation Invariance describes a function that produces identical results regardless of the order of its input. For example, a function that computes the average of the elements in a given array is permutation invariant because the arrangement of elements does not influence the outcome. Figure 2 (b) illustrates this concept using the average operation in the context of matrix multiplication, where $n$ represents the number of rows and $R$ denotes a row-permuted version of the input. As shown, despite reordering the rows, the final result remains unchanged (i.e., $\mathbb{E}[X]=\mathbb{E}[X^{R}]$). Without loss of generality, if a function $f(\cdot)$ satisfies permutation invariance, then for a given input $X$ and a permutation function $\pi(i)$, the lemma $\pi(f(X))=f(\pi(X))$ holds.
It is important to note that matrix multiplication, in general, is not a permutation invariant operation. Specifically, if the weights used in the multiplication are not uniform (e.g., $\frac{1}{n}$ for all elements, as in the averaging operation shown in Figure 2 (b)), the result will differ depending on the order of the rows. This is because such weights can be interpreted as a weighted sum, being sensitive to the input arrangement.
Neural Networks and Shuffling. Modern neural networks exhibit an intriguing property that achieves nearly identical training results even when the arrangement of elements in the input data is shuffled, demonstrating similar behavior to permutation invariance (Tolstikhin et al., 2021; Naseer et al., 2021). Shuffling, here, refers to rearranging the elements of a data sample, such as altering the position of pixels in an image. Similarly, reordering ground truth labels has minimal impact on training outcomes, reflecting characteristics similar to permutation equivariance. However, it is important to note that neural networks exhibit a semi-permutation equivariance and invariance property, as results are not perfectly identical but still closely adhere to these principles. This property arises primarily from the reliance of neural networks on matrix multiplication for most of their operations (Zhang et al., 2019; Kim and Ko, 2022).
Leveraging this property, neural networks can achieve similar training results with spatially shuffled data as with unshuffled data. Since shuffled data is semantically uninterpretable to humans, prior research has highlighted its potential to enhance system privacy. For instance, Amin et al. used permutation equivariance to secure medical image processing (Amin et al., 2024), while Xu et al. explored its applications in transformers for model encryption and privacy-aware split learning (Xu et al., 2024).
3. Feasibility Study and Core Idea
We now present our preliminary studies, which demonstrate the feasibility of permuting data during neural network training from the perspectives of data privacy and model performance preservation. Building on the insights gained from these experiments, we introduce the core concept underlying our proposed framework Tazza.
Figure 3. (a) Bottom-4 (top green box) and Top-4 (bottom red box) PSNR samples from the MNIST dataset. (b) PSNR distribution between the original and shuffled images from the MNIST dataset (using 60K images).
Feasibility Study. To evaluate the feasibility of leveraging shuffling methods for preserving data privacy, we measured the Peak Signal-to-Noise Ratio (PSNR) between original and shuffled samples in the MNIST dataset. For shuffling, we randomly rearranged the pixels in the input data following a consistent and fixed order applied uniformly across all samples. Figure 3 presents sample original/shuffled images and the distribution of measured PSNR values between 60,000 original and shuffled images. A high PSNR indicates structural similarity between two images, whereas a low PSNR suggests significant differences. As Figure 3 (a) shows, the shuffled samples do not retain the semantic information of the original images. Moreover, Figure 3 (b) shows that most samples exhibit low PSNR in the range of 6–8 dB. Even for the top-4 PSNR samples, the shuffled images remain semantically uninterpretable to human observers, confirming the effectiveness of shuffling in preserving data privacy.
Figure 4. Average accuracy of models trained with original data and shuffled data for 10 random seeds.
Next, we trained a 3-layer MLP model for 20 epochs using two dataset configurations: one with original data and the other with shuffled data, as exemplified in Figure 3 (a). For the shuffling process, the test data and training data were shuffled using the same order. Figure 4 presents the average validation accuracy for 10 different random seeds. As the plots show, the model achieves (nearly) identical accuracy regardless of whether the input data were shuffled, demonstrating the neural network’s semi-permutation invariance property. This result suggests that data shuffling can preserve computation consistency in both training and inference, while effectively addressing privacy concerns.
Challenges and Core Idea. As training results remain nearly identical even when using shuffled data, training models on shuffled inputs rather than original can be a plausible approach for preserving data confidentiality. However, while shuffling input data mitigates confidentiality attacks, it still introduces challenges related to computational overhead and system generalizability.
First, models trained on shuffled data require inputs to be shuffled during both training and inference, adding computational overhead for resource-constrained devices in federated learning systems. Second, standard data augmentation techniques, often essential for improving model performance and convergence, rely on spatial or locality-based assumptions (e.g., random cropping or horizontal flipping) (Perez, 2017; Shorten and Khoshgoftaar, 2019). These techniques are incompatible with shuffled data, as shuffling disrupts spatial coherence. Consequently, precomputing shuffled data becomes impractical unless augmentation is specifically adapted to the applied permutation, increasing complexity and reducing generalizability.
To overcome these challenges, we exploit the interchangeability of shuffling data and shuffling weights, a property rooted in the permutation invariance of neural networks. Neural network weights trained on shuffled data inherently adapt to the shuffling rules of the input. By carefully rearranging these weights and leveraging the properties of permutation invariance and equivariance, the model can behave as if it were trained on unshuffled (original) data. Conversely, shuffling the weights of a model trained on unshuffled data enables it to operate equivalently with shuffled inputs, thereby preserving the confidentiality of the system.
This weight shuffling approach can alleviate the computational overhead by eliminating the need for repeated data shuffling operations during training and inference while maintaining compatibility with the original data. Additionally, this approach restores the applicability of standard data augmentation techniques, preserving the framework’s generalizability and reducing complexity.
4. Threat Model
This work considers two key stakeholders in federated learning scenarios: clients and the server. Both clients and the server can potentially be victims or act as adversaries. Thus, we address two different types of attackers namely integrity attacker, and confidentiality attacker.
Goals and Capability First, an integrity attacker refers to a malicious client that aims to compromise the performance of the global model by injecting harmful information through tampered model updates. Integrity attackers can launch model or data poisoning attacks, manipulating the local training process by altering labels, training data, or model weights. Second, we consider an honest-but-curious server as a confidentiality attacker, which seeks to infer a client’s training data by exploiting the received information. Given that clients transmit their model updates to the server for aggregation, the attacker has access to the updates explicitly shared by the clients. Furthermore, there can be additional shared information, which can be any type of data that the clients are required to transmit for the global model training process (i.e., amount of training data used).
Assumptions For the integrity attack, we assume that the attacker has no control over its device beyond the local training process. Specifically, the attacker cannot interfere with the weight aggregation, distribution, or client selection processes carried out by the federated learning server, nor can it access information from other benign clients, such as their data distribution, model weights, or updates. Finally, we assume that the proportion of attackers cannot exceed 50% of the total number of clients (Fang et al., 2020), all being common assumptions in prior work. In the case of a confidentiality attacker, the attacker cannot interfere with or access the local operations performed on clients, nor can it eavesdrop on network communications. Lastly, we note that both types of attackers make every effort to achieve their objectives and circumvent defense mechanisms, while the victim systems operate cooperatively to prevent these attacks (Kabir et al., 2024).
Figure 5. Overall workflow of Tazza.
5. Framework Design
5.1. Overview of Tazza
We now detail the design of Tazza, a secure and efficient federated learning framework that simultaneously addresses integrity and confidentiality attacks. Specifically, Tazza features two core modules, weight shuffling and shuffled model validation and exploits the permutation equivariance and invariance properties of neural networks.
Figure 5 illustrates the overall workflow of Tazza: 1 The federated learning server distributes a global model to each client. 2 Clients selected to participate in training share a shuffling rule via a safeguarded peer-to-peer communication. As a communication-efficient approach, Tazza simply shares the random seed for generating the shuffling order (Sec 5.2). 3 Participants of the federated learning round train the neural network model with their locally collected data. 4 Upon completing local training, each participant shuffles its local model weights with respect to the shared rule to ensure the confidentiality of the training operations (Weight Shuffling-Sec. 5.3). 5 Participating clients transmit the shuffled model parameters and a small number of input samples (as small as one) to the server. We note that the input samples are also shuffled to prevent the confidentiality of the system from being compromised. 6 With the updated/shuffled models and data samples, the server performs an integrity check via shuffled model validation on the data samples using each client’s model by measuring the similarity of resulting output vectors (Sec. 5.4). Since models and data samples are shuffled to satisfy the permutation invariance of the neural network, the resulting output vectors are still legitimate. 7 Exploiting the similarity between the clients calculated in the previous step, Tazza employs a clustering algorithm to cluster the clients. As malicious and benign models exhibit different computation results, grouping the clients based on their similarity effectively assures that benign clients are not affected by malicious model updates.
To summarize, in Step 4, the weight shuffling process ensures that by shuffling the weights, the server is not able to infer information about the client’s local data through model analysis. Nevertheless, since weights are shuffled while preserving permutation invariance, the vectors calculated in Step 6 still remain valid. This enables the shuffled validation process to effectively identify malicious clients attempting poisoning attacks by clustering them in Step 7.
5.2. Shuffling Rule Sharing
Before delving into the detailed design of Tazza, we first address a potential issue regarding clients independently shuffling their model weights in differing orders, which would void the validation process and lead to degraded aggregated model performance. Therefore, it is important for clients to agree upon a unified weight-shuffling order. However, the challenge here is that if the server gains access to this unified shuffling rule, it could potentially revert the shuffled models to their original state, posing a security risk. To mitigate this, the exchange of shuffling rules must be conducted exclusively among the clients without server involvement.
Secure peer-to-peer communication protocols, such as zero-knowledge proof (Boo et al., 2021) and Byzantine fault-tolerant methods (Castro et al., 1999), provide practical and well-established mechanisms for achieving this client-only data exchange. As noted earlier, we assume that clients leverage these secure communication methods to share shuffling rules amongst themselves, ensuring that the server remains uninformed of the shuffling specifics. This assumption supports the integrity and confidentiality of the proposed system.
Specifically, Tazza employs a Practical Byzantine Fault Tolerant (PBFT) mechanism (Castro et al., 1999) for exchanging its shuffling rules. The server selects one participating client as the head, which distributes a shuffling rule to others. Participants that receive the rule then broadcast the rule to non-participating clients to reach consensus. Even if a malicious client is chosen as the leader, PBFT ensures that consensus is achieved and the shuffling rule remains secure and consistent across all clients, despite potential attacks.
5.3. Weight Shuffling
The weight shuffling module in Tazza aims to prevent data leakage while preserving computational correctness by leveraging permutation equivariance and invariance. As the name suggests, this operation shuffles the neural network weights of a model updated at the client before sharing them with the server for aggregation. For clarity, we explain the weight shuffling process in the context of matrix multiplication, given its predominant role in neural network computations. It is worth noting that implementation details of this module can vary depending on the baseline federated learning model architecture due to differences in design and computational complexity, as discussed in Section 6.
In essence, weight shuffling in Tazza consists of two core operations: row shuffling and column shuffling. Consider a matrix multiplication between the input matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$ and the weight matrix $\mathbf{W}\in\mathbb{R}^{d\times h}$, where $N$, $d$, and $h$ denote the number of vectors, input dimension, and embedding dimension, respectively. To shuffle the weight rows, a row-permutation matrix $R$ can be applied to $\mathbf{W}$ (i.e., $\mathbf{R}\times\mathbf{W}$). This operation is computationally equivalent to directly shuffling the input, thereby ensuring data confidentiality. Furthermore, by applying the same permutation to the columns of $\mathbf{X}$ via $\mathbf{R}^{\top}$ (i.e., $\mathbf{X}\times\mathbf{R}^{\top}$), the computation results remain unchanged due to the property of permutation invariance.
The weights can be shuffled further with respect to the columns by multiplying the weight matrix $\mathbf{W}$ with a column-permutation matrix $\mathbf{C}$ (i.e., $\mathbf{W}\times\mathbf{C}$). Since column shuffling affects only the column order of the output and not the computation itself, no additional permutation needs to be applied to the input. However, when considering the bias term $\mathbf{b}\in\mathbb{R}^{h}$ in a linear layer (i.e., $\mathbf{X}\times\mathbf{W}+\mathbf{b}$), the same column permutation must be applied to $\mathbf{b}$ to ensure proper alignment with the shuffled weights. Since the output of a layer serves as the input for the subsequent layer, the column shuffling order of the current layer can be directly inherited as the row shuffling order for the weights of the next layer.
This shuffling operation for matrix multiplication (i.e., linear layers) leverages permutation invariance and equivariance, which form the foundation of most neural network computations. However, simply uniformly applying this approach to all weights may not always work as intended. This is because, while the underlying principle remains consistent, specific behaviors and designs of models can vary. Using we elaborate on how this shuffling operation can be adapted for various neural network architectures in Section 6.
5.4. Integrity Check
Integrity check in Tazza is designed to identify and isolate malicious models from benign ones through two key steps: shuffled model validation and cluster-aware aggregation.
5.4.1. Shuffled Model Validation
In the Shuffled Model Validation process, the server utilizes client-provided samples as inputs to the uploaded models. As previously mentioned, these data samples are also shuffled to obscure their semantics, ensuring data confidentiality. Despite this obfuscation, the corresponding weight shuffling maintains computational validity, allowing the server to trust the outputs based on the assumed correctness of the shuffle rules. If an attacker uploads models with mismatched data and weight shuffle orders, inconsistent results will emerge, distinguishing these models from legitimate ones. This forces attackers to comply with the correct shuffle rules to avoid exclusion during subsequent stages.
Malicious models produce outputs that deviate from those of benign models, revealing distinctive patterns. Tazza identifies such deviations by computing pairwise cosine similarity among the output vectors of all models (Park et al., 2023). Malicious models typically exhibit lower cosine similarity with benign ones, providing a basis for their classification and isolation.
Clients are then clustered based on their pairwise cosine similarities, with the similarity matrix ($S$) transformed into a distance matrix ($1-S$) to interpret similarity as a distance metric, which is used as input to the DBSCAN algorithm (Ester et al., 1996). Tazza leverages DBSCAN for its ability to identify outliers and its independence from predefined cluster numbers, making it particularly well-suited for federated learning scenarios where the presence of attackers is unknown.
5.4.2. Cluster-aware Aggregation
Finally, we implement cluster-aware aggregation to fully segregate malicious models from benign ones. In each federated learning round, computed cluster labels serve as pseudo-labels, reflecting the similarities among clients within that round only. These labels are independent across rounds (e.g., labels from round $t$ have no relation to those from round $t+1$). To ensure consistency, Tazza assigns global cluster labels by leveraging clustering histories, enabling effective cross-round clustering. Clients are categorized into two groups: (1) those participating for the first time, who lack a global cluster label, and (2) those with prior participation and an assigned global cluster.
Upon performing shuffled model validation, if all clients in a cluster are first-time participants without any global cluster assignment, Tazza creates and assigns new global clusters. Conversely, if a pseudo-cluster contains clients with pre-existing global cluster labels, all clients in the cluster are assigned to the global cluster label of those clients.
Furthermore, if clients who were previously part of different global clusters are grouped in the same pseudo-cluster during the current round, Tazza merges their global clusters. However, instead of merging all clients from both clusters indiscriminately, Tazza adopts a conservative approach by reassigning only the overlapping clients to the new cluster. This precautionary strategy ensures that a single attacker infiltrating a benign cluster cannot lead to all attackers being included in the same benign cluster.
When global clusters are assigned for participating clients, Tazza performs aggregation within each cluster based on the clustering results, producing an updated model for each cluster. We note that Tazza employs averaging as the aggregation algorithm while alternative methods such as median or trimmed mean can also be applied, offering flexibility in system design. By aggregating models only within the same cluster, Tazza prevents compromised models from contaminating the global model shared across other clusters.
This cluster-specific aggregation ensures that the integrity of the federated learning system is maintained, as malicious models cannot propagate their impact to benign clusters (Liu et al., 2021). Furthermore, by leveraging shuffled model validation, Tazza achieves this robust defense without exposing or interpreting the semantics of clients’ data, thereby preserving both system security and data confidentiality. Overall, this dual protection reinforces Tazza’s effectiveness in safeguarding federated learning processes against security threats.
6. Implementation
As previously noted, details of the weight shuffling operation (and implementations) can vary for different model architectures. In this section, we provide a detailed explanation of the weight shuffling process for representative architectures.
Multi-layer Perceptron (MLP). As a simple architecture, we demonstrate the weight shuffling operation in an MLP. In an MLP with $L$ layers, the output $y$ is computed as follows:
(1) $\begin{split}h^{(0)}&=x,\\ h^{(l)}&=\sigma(W^{(l)}h^{(l-1)}+b^{(l)}),\quad l=1,\dots,L-1,\\ y&=W^{(L)}h^{(L-1)}+b^{(L)},\end{split}$
Here, $h^{(l)}$ denotes the activations of the $l$-th layer, $W^{(l)}$ represents the weight matrix associated with the $l$-th layer, $b^{(l)}$ is the bias vector for the $l$-th layer, and $\sigma(\cdot)$ is a non-linear activation function such as ReLU, sigmoid, or tanh.
A critical observation is that commonly used activation functions are parameter-free and permutation-equivariant, meaning their intrinsic behavior remains unaffected by input shuffling. As a result, activation functions can be excluded from considerations in weight shuffling operations. This leaves the linear transformations in each layer, represented as matrix multiplications ($W^{(l)}h^{(l-1)}$), as the primary focus. The weight matrices and bias vectors in these transformations can be shuffled according to the previously outlined strategies, ensuring that computational outcomes remain consistent despite weight shuffling. By applying this shuffling mechanism iteratively across layers, the entire MLP architecture retains compatibility with weight shuffling operations, preserving computational integrity while safeguarding model confidentiality.
Figure 6. Workflow of weight shuffling on a sequential model. For simplicity, we omit the bias addition.
Sequential Models (RNN and LSTM). For recurrent sequential models such as RNNs and LSTMs, the weight shuffling process is similar to that of MLPs, with one key distinction: the inclusion of hidden state operations. These operations enable sequential models to capture and process temporal dependencies in input data. Figure 6 illustrates the fundamental operation of a recurrent model (e.g., RNN), where the next output is computed by integrating the current input ($x_{t}$, $W_{ih}$) with the hidden state vector from the previous time step ($h_{t-1}$, $W_{hh}$). Despite this added complexity, the core computations in sequential models are rooted in matrix multiplications. This ensures that they, like MLPs, exhibit permutation equivariance and invariance properties, allowing weight shuffling to be applied without compromising computational correctness.
Notably, this property extends beyond RNNs to other popular sequential architectures, such as LSTMs, Bi-LSTMs, and GRUs. These models rely on matrix operations for processing both inputs and hidden states, making them equally compatible with weight shuffling techniques.
Figure 7. Workflow with a transformer baseline. For clarity, we omit the bias addition and normalization.
Transformers. Transformers, a widely used modern neural network architecture, operate in three main steps, as depicted in Figure 7. First, the input is segmented into a sequence of tokens, which are passed through a linear embedding layer to generate token embeddings. These embeddings are then processed by multiple encoder layers that analyze the input. Each encoder employs a multi-head attention mechanism comprising three learnable matrices: $W_{Q}$ (query), $W_{K}$ (key), and $W_{V}$ (value). Finally, the encoded feature representation, $Z$, is fed into the classification head (i.e., CLS Head) to produce the final output. Importantly, the encoded feature $Z$ preserves the order of the input tokens, ensuring the model’s output aligns with the input sequence.
From a data perspective, shuffling can be applied at two levels: intra-token and inter-token. At the intra-token level, individual elements within a token are permuted. At the inter-token level, the sequence order of tokens is rearranged. Importantly, since all tokens pass through a linear embedding layer—essentially a matrix multiplication operation—the permutation invariance property of matrix multiplication ensures that identical computational results can be achieved by shuffling the embedding weights accordingly.
For inter-token shuffling, the sequential order of tokens is initially preserved during the attention mechanism (e.g., row order $e_{1},e_{2},\dots,e_{n}$), where spatial or sequential positions are learned and encoded. However, in the final CLS Head, these positional features determine the model’s predictions. Consequently, inter-token shuffling can be achieved by appropriately shuffling the weights of the classification head, ensuring consistent results while altering the token order.
We note that operations beyond the matrix multiplications discussed above (e.g., layer normalization, softmax) exhibit permutation equivariance (e.g., the softmax function) or permutation invariance (e.g., mean and variance computations in layer normalization). Consequently, these operations are unaffected by weight shuffling, ensuring that the computational results remain consistent. This allows Tazza to uphold not only the confidentiality of client data but also the integrity of the system.
7. Evaluation
Now we evaluate Tazza with thorough and extensive experiments with four datasets and various comparison baselines.
7.1. Experiment Setup
Dataset and Model. We evaluate Tazza using four datasets with corresponding model architectures suitable for mobile federated learning scenarios as we detail below.
$\bullet$ MNIST (Deng, 2012) dataset consists of 60,000 training and 10,000 test grayscale handwritten digit images with 28$\times$28 pixel resolution. We use a 3-layered MLP (Rosenblatt, 1958) for evaluating MNIST.
$\bullet$ CIFAR10 (Krizhevsky et al., 2009) is a benchmark dataset consisting of 60,000 32$\times$32 pixel color images , categorized into 10 distinct classes. For this dataset, we use the MLP-Mixer (Tolstikhin et al., 2021) as the model, a design similar to the Vision Transformer but substituting self-attention blocks with MLP-based layers.
$\bullet$ STL10 (Coates et al., 2011) consists of 13,000 natural images across 10 classes, with a resolution of 96$\times$96 pixels. Its higher resolution makes it ideal for assessing Tazza’s scalability with image resolution. For this dataset, we employ Vision Transformer (Dosovitskiy et al., 2020) to manage the increased complexity.
$\bullet$ MIT-BIH Arrhythmia Database (Moody and Mark, 2001, 1992) is a widely used electrocardiogram (ECG) benchmark dataset comprising approximately 70K five-second ECG signals. We implement an RNN  (Rumelhart et al., 1986) to classify three arrhythmia types along with normal cardiac patterns. This dataset is utilized to explore the potential extension of Tazza to time-series inputs.
Baselines. For confidentiality attacks, we evaluate three baseline defense mechanisms: no defense, gradient pruning (Zhu et al., 2019), and differential privacy (Abadi et al., 2016; Dwork et al., 2014). Gradient pruning zeros out small-amplitude gradients to reduce data leakage but requires a carefully balanced pruning ratio; higher ratios mitigate leakage but risk discarding useful knowledge, impacting both performance and convergence. Differential privacy, implemented via DPSGD (Abadi et al., 2016), adds perturbation to gradients to obscure sensitive information, with perturbation levels divided into small, medium, and large, as in prior work (Abadi et al., 2016). Similar to gradient pruning, differential privacy also requires careful tuning of perturbation levels to balance between accuracy and privacy.
For integrity attacks, we evaluate two baseline configurations: FedAvg (McMahan et al., 2017) and Oracle, representing naive and ideal model aggregation scenarios, respectively. Specifically, FedAvg does not employ security mechanisms, leaving it vulnerable to malicious updates, while Oracle assumes perfect knowledge of attackers, with the capability of fully excluding malicious models’ contributions on the global model.
We also evaluate robust aggregation methods leveraging statistical priors, including Median and Trimmed Mean (Yin et al., 2018). Median replaces averaging with the median to mitigate extreme outlier influence, while Trimmed Mean excludes outliers before averaging, reducing the impact of malicious updates. Additionally, we compare Tazza with MultiKrum (Blanchard et al., 2017) and FLTrust (Cao et al., 2020). MultiKrum selects updates with the smallest pairwise Euclidean distances after discarding the most distant ones. FLTrust assigns trust scores based on cosine similarity with a server-trained reference model, adjusting aggregation weights to suppress malicious updates.
Attack Methods. For confidentiality attacks, we use the inverting gradient attack (Geiping et al., 2020) for 100 test samples as a baseline. This method reconstructs training data by optimizing inputs to produce gradients similar to those sent by the client, achieving state-of-the-art performance with minimal additional assumptions (Huang et al., 2021).
For integrity attacks, we evaluate three poisoning attack types: label flipping, noise injection, and backdoor attacks (Fang et al., 2020). In label flipping, the labels of a subset of training data are altered to mislead the locally trained model. Noise injection introduces random noise into the model parameters before they are sent to the server, aiming to degrade the global model’s performance. Backdoor attacks embed specific triggers into the training data, causing the model to behave maliciously when exposed to those triggers.
Miscellaneous Configurations. Throughout the evaluation, unless otherwise specified, datasets are distributed across 100 clients, consisting of 25 attackers, with data shared via a Dirichlet distribution with $\alpha$=0.5 (Hsu et al., 2019). In each federated learning round, 25 participants are selected, and their models are trained locally for 5 epochs using the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 1e-3 and a batch size of 64. Besides the evaluations performed on real embedded computing platforms in Section 7.5, all experiments were conducted on a server with an NVIDIA RTX 2080 GPU, an Intel Xeon Silver 4210 CPU 2.20GHz, and 64GB RAM.
7.2. Overall Performance
Figure 8. LPIPS and accuracy of varying federated learning configurations. $\blacktriangle$ and $\bullet$ denote the applying of gradient pruning and DPSGD, respectively.
To evaluate the overall performance of Tazza, we conducted experiments using the MNIST-MLP and CIFAR10-Mixer configurations with 75 benign clients and 25 label-flipping attackers (i.e., integrity attack), totaling 100 clients. To understand Tazza’s confidentiality-preserving capabilities, the server executed an inverting gradient attack. The effectiveness of the attack was measured using Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), which quantifies semantic distance in image data. Lower LPIPS values indicate greater similarity between the original training samples and those reconstructed by the attack, signifying higher attack success.
Figure 8 presents the global model accuracy under label-flipping attacks and the LPIPS across various integrity and confidentiality defense schemes. The legend indicates different integrity defense schemes, with triangle ($\blacktriangle$) and circle ($\bullet$) markers representing confidentiality defense schemes using gradient pruning and DPSGD, respectively. Gradient pruning is applied with pruning ratios of 70%, 90%, and 99%, while DPSGD uses three perturbation levels: small, medium, and large, as outlined in prior work (Abadi et al., 2016). Each scheme includes three plots: the leftmost plot corresponds to the weakest confidentiality defense configuration (70% pruning for gradient pruning and small perturbation for DPSGD), while the rightmost plot represents the strongest configuration (99% pruning for gradient pruning and large perturbation for DPSGD).
Existing confidentiality defense mechanisms improve data privacy, achieving higher LPIPS values, but often result in reduced accuracy. In contrast, Tazza achieves both high LPIPS and high accuracy, demonstrating its robustness against integrity and confidentiality attacks. Notably, even under an idealized (Oracle) scenario where the server perfectly excludes malicious updates during aggregation, the model accuracy remains lower than that of Tazza, due to inaccuracies introduced by confidentiality defense mechanisms in the aggregated weights.
7.3. Confidentiality Attack Mitigation
Figure 9. Sample confidentiality attack results for different dataset/defense configurations.
FedAvg GP 0.9 DP-M Tazza Avg PSNR ($\downarrow$) MNIST 47.08±12.33 4.92±0.53 6.88±2.17 4.95±0.21 CIFAR10 18.11±6.77 13.17±3.73 15.60±4.01 11.74±2.92 STL10 12.21±4.88 10.23±2.33 12.14±4.67 9.91±2.38 MITBIH 23.56±11.09 -7.37±3.78 19.33±13.39 17.01±8.74 Best PSNR ($\downarrow$) MNIST 61.28 6.34 14.77 5.31 CIFAR10 30.46 20.47 24.33 21.12 STL10 21.32 15.15 20.08 16.18 MITBIH 35.35 0.08 32.34 22.69 Avg LPIPS ($\uparrow$) MNIST 0.04±0.16 0.69±0.04 0.52±0.09 0.75±0.02 CIFAR10 0.36±0.24 0.61±0.10 0.51±0.12 0.62±0.07 STL10 0.63±0.17 0.73±0.07 0.68±0.12 0.75±0.05 Best LPIPS ($\uparrow$) MNIST 0.00 0.54 0.33 0.69 CIFAR10 0.07 0.37 0.31 0.40 STL10 0.34 0.62 0.50 0.58
Table 1. Average and highest PSNR/LPIPS between original and inferred sample results from confidentiality attack for different dataset/defense configurations.
To evaluate the efficacy of Tazza in mitigating confidentiality attacks, we measured PSNR and LPIPS between the original training samples and those inferred from the attack. PSNR assesses structural similarity, while LPIPS evaluates semantic distance for image data. Note that, as LPIPS is specific to images, only PSNR is reported for the MIT-BIH time-series dataset. Table 1 presents the average and best PSNR and LPIPS for various dataset and defense scheme configurations. The best values represent the most similar samples among the 100 used in the experiments. Higher PSNR and lower LPIPS indicate greater similarity between inferred and original samples.
As shown in Table 1, FedAvg without defense mechanisms results in high PSNR and low LPIPS, indicating substantial local data leakage. In contrast, defenses such as gradient pruning with 90% pruning (GP 0.9), DPSGD with medium perturbation (DP-M), and Tazza demonstrate competitive reductions in PSNR and increases in LPIPS, effectively mitigating confidentiality attacks. Notably, while gradient pruning and differential privacy face challenges in balancing model accuracy with confidentiality, Tazza successfully defends against attacks without compromising accuracy.
[…]
7.4. Integrity Attack Mitigation
Figure 10. (a) Global model accuracy and backdoor accuracy for different defense schemes against backdoor attack. (b) LPIPS and accuracy of different defense mechanisms under noise injection attack.
In this section, we evaluate Tazza’s robustness against various integrity attacks, specifically focusing on noise addition and backdoor attack scenarios. The experiments exploit the CIFAR10-Mixer configuration with medium-level DPSGD as the confidentiality defense mechanism for all baselines. For the backdoor attack, attackers assign specific labels to samples by injecting a 4$\times$4 red square patch into the upper-left corner of 50% of their training data as the backdoor trigger (Lee et al., 2024). For the noise addition attack, Gaussian noise with a scaling factor of 0.25 is injected into the model weights before they are uploaded to the server.
Figure 10 presents the global model accuracy (higher values indicate better performance on normal samples; y-axis) and backdoor accuracy (lower values indicate better attack mitigation for triggered samples; x-axis). As shown, baseline methods struggle to effectively counter backdoor attacks and often sacrifice global accuracy. On the other hand, Tazza achieves an 18.89% reduction in backdoor accuracy compared to the FedAvg baseline while maintaining a global model accuracy of 55.37%.
Figure 11. Visualization of client clustering. Different colors denote different clusters. Best viewed in color.
To effectively show how benign and malicious clusters are separated in Tazza, we devise another experiment with 10 benign clients (Clients 1-10) and 10 backdoor attackers (Clients 11-20) training the MNIST-MLP configurations for 50 federated learning rounds. We visualize the cluster allocation for each client at each training round in Figure 11. The illustration clearly shows that Tazza effectively isolates benign clients from malicious attackers demonstrating the efficacy of Tazza in mitigating integrity attacks.
7.5. Scalability on Real Embedded Platforms
Figure 12. Latency of different confidentiality attack mitigation schemes with varying (a) hidden layer counts and (b) number of local training updates.
To evaluate the scalability of Tazza and suitability for real-world applications, we conducted experiments on three embedded computing platforms: Raspberry Pi 3B (Pi, 2023), Nvidia Jetson Nano (Nvidia, 2019a), and Nvidia Jetson TX2 (Nvidia, 2019b). These experiments focus on the computational latency of executing Tazza compared to widely used attack mitigation schemes.
Figure 12 plots the computational latency of various defense mechanisms across different network depths (Figure 12 (a)) and numbers of mini-batch updates during local training (Figure 12 (b)) using the MNIST-MLP configuration on the Raspberry Pi platform. In Figure 12 (a), latency increases with neural network depth, exhibiting a near-linear trend across all methods. However, Tazza demonstrates a less-steep increase compared to alternative defense schemes such as gradient pruning and differential privacy. Similarly, Figure 12 (b) shows a similar trend with the number of mini-batch updates, where Tazza’s advantage becomes even more pronounced. This efficiency comes from the fact that approaches such as gradient pruning and differential privacy require additional operations at every training epoch, whereas Tazza performs only a single shuffling operation once training is complete.
Figure 13. Latency of different confidentiality attack mitigation schemes with varying model architectures
In Figure 13, we focus on the latency of operations in each scheme, excluding the time spent on model training to isolate the overhead introduced specifically for confidentiality attack mitigation. This experiment, conducted on a Raspberry Pi 3B, evaluates various model architectures to account for differences in computational overhead across models. The results in Figure 13 show that Tazza achieves the most efficient performance, primarily due to its reliance on a single shuffling operation after training completes, minimizing additional computational overhead.
Figure 14. Latency of varying confidentiality attack mitigation schemes for different embedded platforms.
Finally, Figure 14 depicts the latency of various defense schemes operating on different embedded platforms using an MLP model with a single hidden layer and 10 mini-batch updates. Overall, we can observe that Tazza shows superior performance compared to alternative mitigation approaches, by Tazza achieving up to 6.70$\times$ lower latency compared to gradient pruning on the Raspberry Pi 3B. Furthermore, as a positive side-effect, Figure 14 shows that Tazza effectively reduces computational disparities among heterogeneous client platforms, with a 2.01$\times$ difference in latency between the most powerful and weakest devices, compared to 2.68$\times$ for gradient pruning. These results serve as evidence that Tazza is suitable for mobile/embedded federated learning systems.
8. Discussion
$\bullet$ Permutation Variant Neural Networks. This work proposes Tazza, a framework addressing two core security challenges in federated learning by leveraging the permutation equivariance and invariance properties of neural networks. While these properties are commonly observed in modern neural networks, certain architectures (e.g., CNN) that lack these properties still persist (Battaglia et al., 2018). While they can often be replaced with alternatives that exhibit these properties, there are cases where the model best suited for a specific application does not qualify this criteria. As a result, these models cannot directly leverage the weight shuffling mechanism in Tazza. However, it is important to note that many of these neural networks are optimized at the software level and are eventually transformed into matrix multiplication operations (Chellapilla et al., 2006; Vasudevan et al., 2017). Considering such optimization, it is feasible to devise new shuffling strategies tailored to these specific models, enabling their integration into the Tazza framework.
$\bullet$ Counter Attack Against Tazza. A key function for Tazza to function effectively is ensuring confidentiality by preventing the server from knowing the shuffling rules. However, the server could attempt counter-attacks to infer this information. For instance, there have been attempts to train models capable of solving jigsaw puzzles to recover original inputs from shuffled data (Li et al., 2021). While such studies have made progress in identifying patch-level shuffling, research in this domain remains incomplete (Carlucci et al., 2019). Furthermore, the server might try to infer shuffling rules through neural network parameter analysis. Nevertheless, as discussed earlier, leveraging the orthogonality of frameworks provides a promising direction to counter such attacks. While an intriguing topic, the primary goal of this study is to demonstrate the feasibility of addressing the two critical security issues in federated learning by exploiting permutation invariance and equivariance. As such, we defer this exploration to future work.
$\bullet$ Network Scalability. From a systems perspective, a critical consideration for Tazza is its scalability. Tazza relies on secure peer-to-peer communication for exchanging shuffling rules. In scenarios where geographically dispersed clients are selected, clients have limited communication capabilities, or the number of clients becomes very large, thus the communication stability and computational or time cost of communication may pose challenges. However, rather than directly sharing full shuffling matrices, a more lightweight approach, such as exchanging random seeds, can significantly reduce the overhead. Nevertheless, a deeper analysis on scalability remains a valuable direction for future work, particularly to address potential bottlenecks and ensure the system’s robustness in more demanding settings.
9. Conclusion
In this paper, we introduce Tazza, a secure, privacy-preserving, and efficient federated learning framework designed to address the intertwined challenges of integrity and confidentiality in decentralized systems. Leveraging the permutation equivariance and invariance properties inherent in modern neural networks, Tazza employs innovative components such as Weight Shuffling and Shuffled Model Validation to enhance resilience against a wide range of attacks, including model poisoning and data leakage. Comprehensive evaluations across diverse datasets and embedded computing platforms demonstrate that Tazza effectively mitigates both integrity and confidentiality threats while maintaining high model accuracy and achieving up to 6.7$\times$ computational efficiency compared to alternative defense schemes. These results highlight the potential of Tazza as a practical and robust solution for secure federated learning in real-world scenarios, catalyzing future research in balancing privacy, security, and performance in decentralized systems.